# Exploring Channel Distinguishability in Local Neighborhoods of the Model Space in Quantum Neural Networks

### v1

@@ 5 Experimental Setup
To account for favorable parameter initializations, we train each architecture 50 times. We use 150 iterations, and monitor convergence and parameter changes. Further, we calculate the channel sensitivity between the quantum channels in every iteration, to get an operationally meaningful analysis of how much the model changes. This also allows evaluating whether the model changes more when updated based on the loss function, rather than by just randomly permuting parameters.
@@
Today’s QNNs use shallow circuits with a depth in $O(\text{log}\ n)$ and a local cost function, as they can be shown to be trainable (Cerezo et al., 2021b). Even if favorable loss landscapes were found for deeper circuits, adding layers results in noisier loss estimates, prohibiting effective learning. Moreover, scaling in number of qubits is limited due to quantum hardware limitations, or, when the models are simulated classically, to the curse of dimensionality. Therefore, many recent works proposing QML to solve a particular problem use few qubits, e.g., Blance and Spannowsky (2021), with two qubits, or Yano et al. (2020) with three and four qubits.
Further, our main goal is to take initial steps in quantifying and characterizing QNN model behavior. As has been mentioned before, due to the power of classical ML, we are unlikely to find an advantage in QML in the near term (Schuld and Killoran, 2022). Therefore, it would be important to focus on more fundamental questions, such as “how can we exploit quantum mechanics for ML purposes?”, to explore the potential of quantum computing for data processing. In that sense, understanding how quantum channels change upon parameter updates is a foundational topic in ansatz design.
@@
Our experiments are implemented using the Python programming language. We use Pennylane (Bergholm et al., 2022) for obtaining the unitary transformations and QuTIP (Johansson et al., 2012, 2013) for calculating the diamond norm. Unfortunately, a diamond norm computation between two operators $\|A-B\|_{\Diamond}$ with a finitely-large overlap $A^{\dagger}B\approx\mathbb{I}$, yields numerical instabilities with the QuTIP package. To account for these numerical issues, the operator can be multiplied with a global phase, which fixes a gauge (cf. Gauge-Fixing, in Appendix D).
6 Results
In the following, we first provide an upper bound on the channel sensitivity based on the applied perturbation, then we discuss our numerical results.
6.1 A Bound on the Channel Sensitivity
We can prove an upper bound on the channel sensitivity through Taylor expansion, which is shown in Equation 7. Therefore, we establish a direct dependence of the distinguishability of the channels to the sum of changes applied [^6] For the precise mathematical derivation, we refer to Appendix C.. The bound holds if the Hermitian generators of the trainable gates are unitary as well, which is relevant for HEAs, as the trainable gates are exponentiations of the $X$, $Y$, and $Z$ Pauli gates. The precision of the bound depends on the magnitude of $\bm{\delta}$, however, considering that the models are trained iteratively in small steps, we expect the bound to hold.
$\begin{split}\|U(\bm{\vartheta})-U(\bm{\vartheta}+\bm{\delta})\|_{\Diamond}&=\|U(\bm{\vartheta})-(U(\bm{\vartheta})+\\ &\sum_{j=1}^{dim(\bm{\delta})}\delta_{j}\frac{\partial U(\bm{\vartheta})}{\partial\vartheta_{j}}+O(\bm{\delta}^{2}))\|_{\Diamond}\\ &\approx\|-\sum_{j=1}^{dim(\bm{\delta})}\delta_{j}\frac{\partial U(\bm{\vartheta})}{\partial\vartheta_{j}}\|_{\Diamond}\\ &\leq\frac{\sum_{j=1}^{dim(\bm{\delta})}|\delta_{j}|}{2}\end{split}$ (7)
6.2 Experiments
@@
We run extensive numerical experiments to validate our bounds. First, we randomly perturb parameters, then we compare them to the channel sensitivity we observe during actual training runs.
Initially, we verify the assumptions about the magnitude of parameter changes. In training runs, the mean parameter change per update in the first 10 iterations is maximum $1.4$%, with a mean and standard deviation of $0.8\%\pm 0.4\%$, but it quickly decreases. We visualize this in Figure 1(a), with the iterations on the x-axis, and the mean change in parameters on the y-axis. Architectures with the same number of qubits are aggregated and show very narrow confidence intervals. Further, it is also visible that changes in multi-qubit systems tend to be even smaller (less than two orders of magnitude). Moreover, we observe that almost all parameters are updated in every iteration, which is shown in Figure 1(b), where the y-axis shows the percentage of parameters changed.
6.2.1 Random Perturbation
@@
The experiments show the same patterns as the bounds predict. In particular, fixing the qubits and parameters per layer while increasing the depth of the circuit, results in a bigger channel sensitivity. The same can be observed when fixing the number of layers and increasing the qubits. We observe that increasing channel sensitivity is more prevalent for larger systems than for deeper circuits, e.g, architectures with 2 qubits, 5 layers and 3 parameters have a smaller channel sensitivity than architectures with 3 qubits, 5 layers and 2 parameters, despite having the same number of parameters.
The behaviour is expected for ansatzes covering large parts of the search space, as the Hilbert space of a larger system is significantly bigger, scaling as $2^{n}$. This allows more indepedent search directions, therefore, potentially more changes in channel upon update. Deeper circuits can explore larger parts of the Hilbert space too, however, their expressivity stagnates when achieving overparameterization (Larocca et al., 2023). In this regime, more parameters will not result in more independent search directions, hence, the difference in channel sensitivity is expected to be smaller.
@@
6.2.2 Training Updates
Additionally, we want to consider how different the quantum channels are when updated based on loss and optimizer. In particular, it could be possible that changing the parameters during training, changes the underlying channel a lot more than perturbing in random directions. Upon analysis of our results, we can confirm the validity of our Taylor expansion even in early stages of iteration, i.e., in the $45,500$ models we trained, the bound holds for every single parameter update taken.
We plot the bounds and channel sensitivity against the iteration in Figure 3 for different layers, qubits and parameters per layer. We mainly observe three things; firstly, the confidence intervals are very narrow, i.e., even though we collect data from 50 different runs for every model, the observable channel sensitivity hardly varies. Further, there is a large discrepancy between our bound and the channel sensitivity, which grows in the number of qubits. Together with our third observation, that the channel sensitivity does not scale as clearly as the bound in the number of parameters, this provides another perspective on training difficulties in today’s QNNs.
While it is not possible to observe the magnitude of channel sensitivity that the bound predicts, the pattern can be observed when looking closely, which is visualized in Figure 4(a), for early training phases, and in Figure 4(b) for later ones. We see that the pattern exists for larger parameter changes in the beginning, however, in fine-tuning or convergence phases, the pattern is not as clear anymore.
(a) Iterations $<$ 15
(b) Iterations $>$ 60
Figure 4: Channel sensitivity scaling in different training phases
@@ 7 Discussion
Our bound establishes a direct dependence of the distinguishability of the channels during QNN training to the number of parameters. That is, the maximum distinguishability of two output states of the quantum channels, scales with (1) the magnitude of changes and (2) the number of parameters. As iterative updates with a small learning rates are applied, which is independent of the number of parameters in the circuit, the dependence is largely dominated by the number of parameters.
This could significantly contribute to the trainability issues of today’s QNNs. In particular, for reasonably sized models, the channels may be distinguishable in early training stages, but hardly distinguishable when fine-tuning the ansatz later. Comparison of our bound to the channel sensitivity obtained in training runs reveals that during training, the channels are even less distinguishable, i.e., our bound seems to be quite loose. While the channel sensitivity for random permutation is still considerably lower than our bounds would predict, a lot more variation can be observed.
This confirms our initial motivation of studying the model space independently of the loss and data, as these are global properties of ansatzes. In particular, we want to draw attention to the remarkable similarities to the BP phenomenon. It seems that the architectures significantly contribute to the untrainable models we observe today, independently of data and loss. Our findings have implications to start a more thorough investigation of the model space of the ansatzes that are currently used, and, ultimately, stress the need for good initialization and warm-starting. It is, however, non-obvious how to find such starting points without trial-and-error.
@@ 8 Conclusion
Considering that QNNs were designed for NISQ technology, thus limited to shallow, small-scale circuits due to hardware limitations, the community will be limited to such small instances in the near-term. Therefore, our findings have direct implications on the meaningfulness of iteratively optimizing such circuits. While it is true that the models can be scaled to work for larger circuits, we have overwhelming evidence through works on BPs, that this is inherently difficult. Together with results on classical simulability and dequantizability in the absence of BPs, this work extends the evidence that we need a paradigm shift in Variational Quantum Computing or even QML altogether.
Reproducibility Statement
We provide the source code for our experiments in a GitHub repository [^7] https://github.com/sabrinaherbst/qnn_local_neighborhood with extensive documentation and listed dependencies.
@@ A.1 Quantum Computing
A qubit is characterized by 4 numbers (the real and imaginary parts of $\alpha$ and $\beta$), which, when translated into polar coordinates, gives a convenient geometrical representation on the so-called Bloch sphere. This plays a significant part in today’s QNNs, as the trainable gates are rotations of the individual qubits around the x-, y- and z-axis of this sphere.
@@ A.2 Inner-product & overlap between vectors
The inner-product is a generalization of the dot-product between vectors and measures the overlap or coherence between vectors sampled from a vector space.
@@ A.3 Unitary groups and the Haar Measure
(Hoggar, 1982) Let $\mathbb{C}^{d}$ denote a d-dimensional Hilbert space with orthonormal basis set {|$\{\ket{n_{k}}\}_{k=1}^{d}$. Owing to their redundancy under global phase transformations and normalization, the quantum states in this space correspond to points in the complex-projective space $\mathbb{CP}^{d-1}$ (Bengtsson and Zyczkowski, 2006). Thus, we define the nontrivial complex-projective t-design as a set of states $S\subsetneq\mathbb{CP}^{d-1}$, sampled according to some probability measure $\mu$, satisfying,
@@ Appendix B Welch bounds and t-designs
(Welch bounds) Let $n\geq d$. If $\big{\{}\bm{v_{i}}\big{\}}_{1\leq i\leq n}$ is a sequence of unit vectors in $\in\mathbb{C}^{d}$, then,
@@ B.1 Simplification of Equation 21a:
$\max_{1\leq j,k\leq n,j\neq k}|\langle\bm{v}_{j}|\bm{v}_{k}\rangle|^{2t}$ $\geq\frac{1}{n-1}\big{[}\frac{n}{\frac{(d+t-1)!}{(d-1)!\ t!}}-1\big{]}$ (23)
$=\frac{1}{n-1}\big{[}\frac{n}{\frac{(d+t-1)(d+t-2)....d}{t!}}-1\big{]}$ (24)
Here in the denominator, we have used the factorial property, $(d+t-1)!/(d-1)!=(d+t-1)(d+t-2)...d(d-1)!/(d-1)!=(d+t-1)(d+t-2)~{}\ldots~{}d$.
@@ C.1 Preliminaries
$\partial_{j}U(\bm{\vartheta})=-iU_{j+1\rightarrow dim(\bm{\vartheta})}H_{j}U_{1\rightarrow j}$ (27)
$\begin{split}U_{l\rightarrow m}(\bm{\vartheta})&=\prod_{j=l}^{m}V_{j}(\vartheta_{j})W_{j}\\ &=\prod_{j=l}^{m}e^{-i\vartheta_{j}H_{j}}W_{j}\end{split}$ (28)
@@ C.2 Derivation of the Bound
We can now analyze the difference in diamond norm upon slight perturbation of parameters in Equation 29. We use a Taylor expansion and truncate after the first order. We arrive at Equation 31 after applying the partial derivative from Equation 27. We use the triangle inequality and homogeneity, which are both necessary conditions for matrix norms (Horn and Johnson, 1985), to arrive at Equation 32 and Equation 33 respectively. Arriving at Equation 34 is non-trivial, as, in general, it is not guaranteed that multiplying hermitian and unitary matrices results in a unitary. However, HEAs use $R_{X}$, $R_{Y}$ and $R_{Z}$ rotations with generators $X$, $Y$ and $Z$ respectively, which are known to be hermitian and unitary. Since unitaries form a group that is closed under multiplication, it forms a valid quantum channel, which evaluates to $1$(Watrous, 2018, Proposition 3.44).
$\|U(\bm{\vartheta})$ $-U(\bm{\vartheta}+\bm{\delta})\|_{\Diamond}$ (29)
$=\|U(\bm{\vartheta})-\bigg{(}U(\bm{\vartheta})+\sum_{j=1}^{dim(\bm{\delta})}\delta_{j}\frac{\partial U(\bm{\vartheta})}{\partial\vartheta_{j}}+O(\bm{\delta}^{2})\bigg{)}\|_{\Diamond}$ (30)
@@
Appendix D Gauge Fixing
Considering the numerical instabilities whenever $A^{\dagger}B\approx\mathbb{I}$, we perform Gauge-Fixing by adjusting the global phases of the two unitaries, which ensures numerical stability without affecting the diamond norm.

### v2

@@ 5 Experimental setup
To account for favorable parameter initializations, we train each architecture 50 times. We use 150 iterations, and monitor convergence and parameter changes. We use the Adam optimizer (Kingma & Ba, 2015) from Pennylane with default parameters (step size: 0.01, beta 1: 0.9, beta 2: 0.99, epsilon: 1e-08). Further, we calculate the channel sensitivity in every iteration, to get an operationally meaningful analysis of how much the model changes. This allows evaluating whether the model changes more when updated based on the loss function, rather than by just randomly permuting parameters. We display aggregated results using confidence intervals with a 95% confidence level.
@@
Today’s QNNs use shallow circuits with a depth in $O(\text{log}\ n)$ and a local cost function, as they can be shown to be trainable (Cerezo et al., 2021b). Even if favorable loss landscapes were found for deeper circuits, adding layers results in noisier loss estimates, prohibiting effective learning. Moreover, scaling in number of qubits is limited due to quantum hardware limitations, or, when the models are simulated classically, to the curse of dimensionality. Therefore, many recent works proposing QML to solve a particular problem use few qubits, e.g., Blance & Spannowsky (2021), with two qubits, or Yano et al. (2020) with three and four qubits.
Further, our main goal is to take initial steps in quantifying and characterizing QNN model behavior. As has been mentioned before, due to the power of classical ML, we are unlikely to find an advantage in QML in the near term (Schuld & Killoran, 2022). Therefore, it would be important to focus on more fundamental questions, such as “how can we exploit quantum mechanics for ML purposes?”, to explore the potential of quantum computing for data processing. In that sense, understanding how quantum channels change upon parameter updates is a foundational topic in ansatz design.
@@
Our experiments are implemented using the Python programming language. We use Pennylane (Bergholm et al., 2022) for obtaining the unitary transformations and QuTIP (Johansson et al., 2012; 2013) for calculating the diamond norm. Unfortunately, a diamond norm computation between two operators $\|A-B\|_{\Diamond}$ with a finitely-large overlap $A^{\dagger}B\approx\mathbb{I}$, yields numerical instabilities with the QuTIP package. To account for these numerical issues, the operator can be multiplied with a global phase, which fixes a gauge (cf. Gauge-Fixing, in Appendix F).
@@
6 Experimental Results
We run extensive numerical experiments to validate our bounds. First, we randomly perturb parameters, then we compare them to the channel sensitivity we observe during actual training runs. Initially, we verify the assumptions about the magnitude of parameter changes. In training runs, the mean parameter change per update in the first 10 iterations is maximum $1.4$%, with a mean and standard deviation of $0.8\%\pm 0.4\%$, but it quickly decreases. We visualize this in Figure 1(a), with the iterations on the x-axis, and the mean change in parameters on the y-axis. Architectures with the same number of qubits are aggregated and show very narrow confidence intervals. Further, it is also visible that changes in multi-qubit systems tend to be even smaller (less than two orders of magnitude). Moreover, we observe that almost all parameters are updated in every iteration, which is shown in Figure 1(b), where the y-axis shows the percentage of parameters changed.
6.1 Random perturbation
@@
The experiments show the same patterns as the bounds predict. In particular, fixing the qubits and parameters per layer while increasing the depth of the circuit results in a bigger channel sensitivity. The same can be observed when fixing the number of layers and increasing the qubits. We observe that increasing channel sensitivity is more prevalent for larger systems than for deeper circuits, e.g, architectures with 2 qubits, 5 layers and 3 parameters have a smaller channel sensitivity than architectures with 3 qubits, 5 layers and 2 parameters, despite having the same number of parameters.
The behavior is expected for ansatzes covering large parts of the search space, as the Hilbert space of a larger system is significantly bigger, scaling as $2^{n}$. This allows more independent search directions, therefore, potentially more changes in channel upon update. Deeper circuits can explore larger parts of the Hilbert space too, however, their expressivity stagnates when achieving overparameterization (Larocca et al., 2023). In this regime, more parameters will not result in more independent search directions, hence, the difference in channel sensitivity is expected to be smaller.
@@
6.2 Training updates
Figure 3: Channel sensitivity during training runs
Additionally, we want to consider how different the quantum channels are when updated based on loss and optimizer. In particular, it could be possible that changing the parameters during training, changes the underlying channel a lot more than perturbing in random directions. Upon analysis of the results, we can confirm the validity of the Taylor expansion even in early stages of iteration, i.e., in the $45,500$ models we trained, the bound holds for every single parameter update taken.
We plot the channel sensitivity in Figure 3 for different layers, qubits and parameters per layer. We can observe that the confidence intervals are very narrow, i.e., even though we collect data from 50 different runs for every model, the observable channel sensitivity hardly varies. Further, the channel sensitivity is substantially smaller than the bound, speaking to the severity of the issue. The discrepancy grows in particular with the number of qubits, which is visualized in Figure 4 in Appendix D. While we cannot observe the magnitude of channel sensitivity that the bound predicts, the pattern of scaling in the number of parameters can be observed. This is in particular visible in Figure 3 for early training stages with larger parameter updates, however, it vanishes in later stages, when the updates are very small. Further, to showcase robustness with respect to the feature encoding (and different parameter updates that may result thereof), we run our experiments with angle encoding as well and verify that the magnitude of channel sensitivity does not change. We refer to Figure 5 in Appendix E for more information.
@@ 7 Discussion
Our bound establishes a direct dependence of the distinguishability of the channels during QNN training to the number of parameters. That is, the maximum distinguishability of two output states of the quantum channels, scales with (1) the magnitude of changes and (2) the number of parameters. As iterative updates with small learning rates are applied, which is independent of the number of parameters in the circuit, the dependence is largely dominated by the number of parameters.
This could significantly contribute to the trainability issues of today’s QNNs. In particular, for reasonably sized models, the channels may be distinguishable in early training stages, but hardly distinguishable when fine-tuning the ansatz later. Comparison of our bound to the channel sensitivity obtained in training runs reveals that during training, the channels are even less distinguishable. While the channel sensitivity for random permutation is still considerably lower than our bounds would predict, a lot more variation can be observed.
This confirms our initial motivation of studying the model space independently of the loss and data, as these are global properties of ansatzes. In particular, we want to draw attention to the remarkable similarities to the BP phenomenon. Adapting a series of composition of maps viewpoint (Larocca et al., 2023), our work argues that the unitaries in the unitary space of the considered architectures do not change significantly upon parameter update in the local neighborhood of the parameter space. Thus, from small channel distinguishability, it follows that the quantum states in the Hilbert space will not differ significantly, as will the expectation values (which is a consequence of the diamond norm formulation), which are directly related to the loss. Roughly speaking, this behavior is mirrored in the BP phenomenon. While we do not argue for strict equivalence, the similarities hint towards two perspectives on similar trainability issues. Therefore, this phenomenon might not only be tied to the exponential Hilbert space, but may be an artifact of the architectures themselves.
Therefore, it seems that the architectures significantly contribute to the untrainable models we observe today, independently of data and loss. Our findings have implications to start a more thorough investigation of the model space of the ansatzes that are currently used, and, ultimately, stress the need for good initialization and warm-starting. It is, however, non-obvious how to find such starting points without trial-and-error.
@@ 8 Conclusion
Considering that QNNs were designed for NISQ technology, shallow, small-scale circuits due to hardware limitations, the community will be limited to such small instances in the near-term. Therefore, our findings have direct implications on the meaningfulness of iteratively optimizing such circuits. While it is true that the models can be scaled to work for larger circuits, we have overwhelming evidence through works on BPs, that this is inherently difficult. Together with results on classical simulability and dequantizability in the absence of BPs, this work extends the evidence that we need a paradigm shift in Variational Quantum Computing or even QML altogether.
@@ Acknowledgments
Reproducibility Statement
The code and results for the experiments are publicly available in a GitHub repository [^8] https://github.com/sabrinaherbst/qnn_local_neighborhood.
@@ A.1 Quantum Computing
A qubit is characterized by four numbers (the real and imaginary parts of $\alpha$ and $\beta$), which, when translated into polar coordinates, gives a convenient geometrical representation on the so-called Bloch sphere. This plays a significant part in today’s QNNs, as the trainable gates are rotations of the individual qubits around the x-, y- and z-axis of this sphere.
@@ A.2 Inner-product & overlap between vectors
The inner-product operation is a generalization of the dot-product between vectors and measures the overlap or coherence between vectors sampled from a vector space.
@@ A.3 Unitary groups and the Haar Measure
(Hoggar, 1982) Let $\mathbb{C}^{d}$ denote a d-dimensional Hilbert space with orthonormal basis set {|$\{\ket{n_{k}}\}_{k=1}^{d}$. Owing to their redundancy under global phase transformations and normalization, the quantum states in this space correspond to points in the complex-projective space $\mathbb{CP}^{d-1}$ (Bengtsson & Zyczkowski, 2006). Thus, we define the nontrivial complex-projective t-design as a set of states $S\subsetneq\mathbb{CP}^{d-1}$, sampled according to some probability measure $\mu$, satisfying,
@@ Appendix B Welch bounds and t-designs
(Welch bounds) Let $n\geq d$. If $\big{\{}\bm{v_{i}}\big{\}}_{1\leq i\leq n}$ is a sequence of unit vectors in $\mathbb{C}^{d}$, then,
@@ B.1 Simplification of Equation 20a:
$\max_{1\leq j,k\leq n,j\neq k}|\langle\bm{v}_{j}|\bm{v}_{k}\rangle|^{2t}\geq\frac{1}{n-1}\big{[}\frac{n}{\frac{(d+t-1)!}{(d-1)!\ t!}}-1\big{]}=\frac{1}{n-1}\big{[}\frac{n}{\frac{(d+t-1)(d+t-2)....d}{t!}}-1\big{]}$ (22)
Here in the denominator, we have used the factorial property, $(d+t-1)!/(d-1)!=(d+t-1)(d+t-2)...d(d-1)!/(d-1)!=(d+t-1)(d+t-2)...d$.
@@ C.1 Preliminaries
$\partial_{j}U(\bm{\vartheta})=-\frac{i}{2}U_{j+1\rightarrow dim(\bm{\vartheta})}H_{j}U_{1\rightarrow j}$ (25)
$\begin{split}U_{l\rightarrow m}(\bm{\vartheta})&=\prod_{j=l}^{m}V_{j}(\vartheta_{j})W_{j}\\ &=\prod_{j=l}^{m}e^{-i\frac{\vartheta_{j}}{2}H_{j}}W_{j}\end{split}$ (26)
@@ C.2 Derivation of the Bound
We can now analyze the difference in diamond norm upon slight perturbation of parameters in Equation 27. We use a Taylor expansion and truncate after the first order. We arrive at Equation 28 after applying the partial derivative from Equation 25. We use the triangle inequality and homogeneity, which are both necessary conditions for matrix norms (Horn & Johnson, 1985), to arrive at Equation 29 and Equation 30 respectively. Arriving at Equation 31 is non-trivial, as, in general, it is not guaranteed that multiplying hermitian and unitary matrices results in a unitary. However, HEAs use $R_{X}$, $R_{Y}$ and $R_{Z}$ rotations with generators $X$, $Y$ and $Z$ respectively, which are known to be hermitian and unitary. Since unitaries form a group that is closed under multiplication, it forms a valid quantum channel, which evaluates to $1$(Watrous, 2018, Proposition 3.44).
$\|U(\bm{\vartheta})-U(\bm{\vartheta}+\bm{\delta})\|_{\Diamond}$ $=\|U(\bm{\vartheta})-\bigg{(}U(\bm{\vartheta})+\sum_{j=1}^{dim(\bm{\delta})}\delta_{j}\frac{\partial U(\bm{\vartheta})}{\partial\vartheta_{j}}+O(\bm{\delta}^{2})\bigg{)}\|_{\Diamond}$ (27)
@@
Appendix D Comparison of empirical channel sensitivity and bound
We compare the empirical channel sensitivity with the predicted bound in Figure 4. We add this figure to visualize that experimentally we can observe a large discrepancy, although the magnitude of this discrepancy makes it hard to read much more from the plot. We also would like to remark that one should be careful in drawing the conclusion that the bound is thus loose, as there might still be a point in the model space where taking such a step could achieve the bound. Our extensive experiments, however, reveal that this will likely not be the case in large areas of the model space.
Appendix E Angle Encoding
We visualize the channel sensitivity when encoding the data with angle embedding. We run these experiments to demonstrate robustness of the channel sensitivity with respect to the changed parameter updates that may occur due the different input states. While the channel sensitivity slightly varies, we can observe that its magnitude does not differ when comparing it to the results using amplitude embedding (we observe a maximum distinguishability of approximately $0.5\%$ in both cases).
Figure 5: Channel sensitivity for angle embedding
Appendix F Gauge Fixing
Considering the numerical instabilities whenever $A^{\dagger}B\approx\mathbb{I}$, we apply a Gauge-Fixing algorithm. In particular, we adjust the global phases of the two unitaries, which ensures numerical stability without affecting the diamond norm.